# EquiTabPFN: A Target-Permutation Equivariant Prior Fitted Network

### v2

@@ 1 Introduction
However, models such as TabPFN [9] are inherently confined to covariate–target pairs of fixed, predefined dimension, thereby limiting their applicability to datasets that match these specifications. This limitation can be alleviated via additional data pre-processing—e.g., projecting high-dimensional covariates into a lower-dimensional subspace—or by post-processing model predictions,—e.g. employing hierarchical strategies for classification tasks with many classes [24, 4]. However, the increased computational burden often offsets some of the benefits offered by in-context learning.
@@ 2 Related Work
Beyond pre-defined target dimensionality. Margeloiu et al. [14] explored training classifiers on features generated by TabPFN. This requires training a model at inference ; unlike our approach. To address TabPFN’s limitation to a predefined number of classes, Hollmann et al. [10] propose to split the classification problem into smaller ones on which the model can be employed, then to aggregate predictions using a strategy, such as the one based on an error-correcting output codes (ECOC) [4]. In this work, we show that such a strategy results in an increased computational cost compared to using our architecture that is natively target equivariant. Qu et al. [22] proposed to use a hierarchical classification strategy [24] which still incurs an increased computational cost. Recently, Wu and Bergman [28] propose a mechanism to handle arbitrary number of classes without target equivariance and require a different paradigm involving an adversarial training procedure.
@@ 3 Background on Prior-Fitted Networks
inline]We probably want to introduce the proj matrix here to be consistent with the input description.
A notable property of the TabPFN architecture is its invariance of the test predictions to the order by which training and test points are provided. This property is desirable since the order of training points is arbitrary and should not influence predictions on test samples. However, the network lacks equivariance w.r.t. the targets’ dimensions, meaning that predictions are highly dependent on the order by which the training target dimensions are provided, an undesirable property as we show in our theoretical analysis in Section 5.
@@ 4 Target equivariant prior-fitted network
We introduce EquiTabPFN, a new model architecture for in-context learning on tabular data, that is permutation equivariant w.r.t. the target’s components. Our architecture integrates self-attention mechanisms across data points and data components to leverage relationships between datapoints while preserving equivariance by processing individual attributes (such as the targets components). Unlike TabPFN which requires fixed target dimensions for all datasets, EquiTabPFN allows the dimensions of the target to change depending on the dataset as a consequence of its equivariant architecture. EquiTabPFN consists of three major modules: a target equivariant encoder, an attention module both across data components and across data points, and a non-parametric equivariant decoder, each designed to facilitate the end-to-end learning of components interactions and datapoint relationships, see Figure 1. Below, we elaborate on each module and their interactions.
@@ 4.1 Target equivariant encoder
The encoder constructs training and test tokens by applying a linear projection to both covariates and targets so that target equivariance is preserved. Specifically, following [9], each training and test covariate vector $x_{n}$ and $x^{\star}_{m}$ is encoded into a single token of dimension $d$ by applying a linear projection matrix $U$ of size $d\times p$. However, instead of adding a linear projection of the training targets to each corresponding training token, as done in the case of TabPFN (see Section 3), we compute a token for each component $(y_{n})_{j}$ of a training target by multiplying them with an embedding vector $V$ of dimension $d$ for all $1\leq j\leq q$. This operation amounts to applying a $1\times 1$ convolution along the components of each target which preserves target equivariance. Since, the validation target $Y^{\star}$ is not provided to the model as input, it is replaced by a trainable prediction token $W_{\text{pred}}$ of dimension $d$ that is repeated $M\times q$ times to form an initial guess $\tilde{Y}^{0}$ of the target. All these embeddings along with prediction tokens are collected to form a single tensor $E$ of shape $(N+M,q+1,d)$, where the blocks $E_{:N,1,:}$, $E_{N:M,1,:}$ correspond to embeddings of $X$ and $X^{\star}$, $E_{:N,1:q,:}$ represents the embedding of $Y$ while $E_{N:M,1:q,:}$ denotes the initial guess $\tilde{Y}^{0}$ obtained using the prediction token. This tensor is then processed by the attention modules as described next. inline]DS: Write matrix dimensions
@@ 4.2 Self-Attention Mechanisms
Self-attention across components
allows interactions among components within each datapoint. It is applied independently per samples to preserve equivariance w.r.t. to the samples. We further employ a masking strategy that we found useful empirically: forcing target tokens to attend only to the covariate token, while allowing the covariate token to attend to all tokens.
Self-Attention across datapoints
captures relationships between datapoint embeddings, allowing the model to aggregate information globally. It is applied between samples and independently per each input dimensions $p$ and $q$ to preserve equivariance. Similarly to [9], training and validation tokens only attend to training tokens.
@@ 4.3 Non-parametric equivariant decoder
The decoder aggregates the processed embeddings to produce prediction $\hat{Y}$. This is achieved in two steps: an attention module first computes an intermediate prediction $\tilde{Y}=(\tilde{y}_{m})_{m=1}^{M}$ in the form of a weighted average of training targets $Y$, then a residual correction is added to produce the final prediction. More precisely, the attention module uses the embeddings of the training and validation samples as keys and queries, while the attention values are simply the training targets $Y$, i.e. $\tilde{y}_{m}=\sum_{n=1}^{N}y_{n}\text{SoftMax}\left(\sum_{i,u}E_{n,i,u}E_{m,i,u}/\sqrt{(1+q)d}\right).$ The residual correction, in the form of a point-wise MLP, operates independently on each dimension $j$ of the attention output $\tilde{y}_{m}$ so that equivariance is preserved while enabling nonlinear interactions between training values.
Without the residual correction and removing the dependence of the keys and queries embeddings on the training targets $Y$ (for instance by setting the weights of the target encoder and pointwise MLP to 00), the decoder becomes a linear non-parametric regression estimator [25, Definition 1.7], which is a generalization of Nadaraya-Watson’s estimator [20, 27]. However, linear estimators are known to be suboptimal compared to non-linear ones [5]. This motivates introducing a nonlinear dependence of the estimator to $Y$, in our setting, to increase the expressiveness of the decoder allowing it to adapt to the prediction task at hand. Experimentally, we also found a clear improvement when adding such a residual correction and making the embeddings dependent on the targets.
@@ 4.4 Pre-training Procedure
Here, $(y,y^{\prime})\mapsto\ell(y,y^{\prime})\in\mathbb{R}$ is a point-wise loss, typically cross-entropy, and the expectation is over the collections datasets sampled according to the dataset prior. Note that each test query $x_{m}^{\star}$ is processed independently by the network $f$ so that $f_{X,Y}\left(X^{\star}\right)=(f_{X,Y}\left(x_{m}^{\star}\right))_{m=1}^{M}$. Next, we show, under natural conditions, that target equivariant functions constitute the right space of functions when searching for solutions to objectives of the form in Equation 1.
@@ 5.1 Optimality of target equivariant networks
It is clear from Definition 5.1 that EquiTabPFN is target equivariant. Even when a model is not target equivariant by construction, it is natural to expect it to learn to be target equivariant, when trained via the objective in Equation 1 over a large class of randomly sampled datasets. To formalize this, we define the target equivariance gap, which quantifies the deviation of a function $f$ from its symmetrized counterpart ${f}^{\text{equi}}$.
@@
Assumption (A) is satisfied for most commonly used losses such as the cross-entropy loss or the quadratic loss. Assumption (B) holds as soon as data can be presented without a preferred ordering, as most i.i.d. tabular datasets. The next proposition, proved in Appendix A, decomposes the objective into a non-negative equivariance gap and an optimality error.
@@
Under Assumptions (A) and (B), the equivariance gap $\mathcal{E}^{\text{equi}}(f)$ is always non-negative and only equal to 00 when $f$ is equivariant to permutations, so that for any $f$:
@@
The key step is to express the objective as an expectation over permutations using Assumptions (A) and (B) on the data and loss: $\mathcal{L}(f)=\mathbb{E}_{p}\mathbb{E}_{\sigma}\left[\ell\left(\sigma^{-1}f_{X,\sigma(Y)}\left(X^{\star}\right),Y^{\star}\right)\right]$. The non-negativity of the equivariance gap is then established using Jensen’s inequality by convexity of the loss $\ell$ in its first argument (Assumption (A)). Now, assume by contradiction that $f^{\star}$ is not equivariant and note that ${f^{\star}}^{\text{equi}}$ is a measurable function by construction. It follows that $\mathcal{E}^{\text{equi}}(f^{\star})>0$, which directly implies that $\mathcal{L}(f^{\star})>\mathcal{L}({f^{\star}}^{\text{equi}})$, thus contradicting the optimality of $f^{\star}$. ∎
Proposition 5.3 shows that minimizing the objective $\mathcal{L}$ results in a target equivariant function. Hence, using a non-equivariant model must employ some of its expressive power solely for being equivariant, which can be wasteful, as we verify empirically in Section 5.2 below.
@@ 5.2 Non-equivariance of TabPFNs models
Predictions instabilities. To illustrate the implications of non-equivariance on the robustness of PFNs models, we consider a toy classification problem in $2$ dimensions, where $9$ training points are positioned on a $2$-dimensional regular grid, each corresponding to a different class inspired by McCarter [15]. Different pretrained models are then used to predict the classes on a regular grid of $40^{2}$ points. Figure 3 shows the classification boundaries when using the same data but with $3$ different orderings for the classes. It is clear that the ordering heavily affects the prediction even in this simple example, which strongly impacts robustness of models like TabPFN and TabPFN-v2. Note that the predictions of TabPFN-v2 are particularly noisy due to having only 9 training data points which is not an issue in itself, unlike the extreme unstability to the class ordering.
Target equivariance gap during training. In Figure 3, we analyse the equivariance error while training TabPFN. Figure 3 (left) shows the equivariance error of TabPFN in terms of percentage of violation of Equation 2, e.g. how frequently the predicted classes $f_{X,\sigma(Y)}(X^{*})$ and $\sigma\left(f_{X,Y}(X^{*})\right)$ differ. We sample 512 datasets from the prior and report standard deviation. In Figure 3 (left), the equivariance error is clear and slowly decreases during training. This non-equivariance 1) induces additional errors for the model as demonstrated in Proposition 5.3 and 2) causes the model to provide surprising predictions given that permuting the output order can change the results as seen in Figure 3.
Mitigation via costly ensembling. Hollmann et al. [9] introduce a mitigation strategy to non-equivariance by averaging several predictions using random permutations of the target dimensions. Averaging over all possible permutations gives an equivariant function as discussed in Equation 4. However, this requires making $\mathcal{O}(q!)$ calls to the model, where $q$ is the number of classes. This becomes quickly prohibitive, even for $q=10$ as considered in the original study. We illustrate how fast the model becomes equivariant when using ensembles in Figure 3 (right). While ensembling more permutation helps to make the model more equivariant, many ensembles are required, in particular, when considering more classes. In contrast, the model we propose is fully target equivariant so that predictions remain the same regardless of the class ordering as illustrated in Figure 3 (bottom).
@@ 6.1 Experimental setup
We trained our model, EquiTabPFN on artificial dataset extending the public code of Müller et al. [17] and using the same procedure from Hollmann et al. [9] employed to train TabPFNv1. Crucially, it was trained on classification tasks with less than $10$ classes. The total training time took approximately 4 days on a single A100 GPU with 80GB of memory. We set our architectures hyperparameters so that they match the number of parameters of previous work, in particular the model we trained contains $\sim$25M parameters similar to the baselines we consider and we match most our hyperparameters to the same values as Hollmann et al. [9]. We refer to Appendix B for further training details and description of the hyperparameters used.
@@
For evaluation, we consider classification tasks from the TabZilla benchmark [16]. To assess the impact of unseen class counts during evaluation (i.e., exceeding $10$), we consider two setups: one on $76$ multi-class tasks with at most $10$ classes, and another on $10$ tasks with more than $10$ classes. Details of the datasets/tasks used are provided in Tables 1 and 2. In all results, except those of Figure 5 (left), we report the results of EquiTabPFN without ensembling as it is not critical for the performance of our method.
@@
6.2 Results
[…]

### v4

@@ 1 Introduction
However, models such as TabPFN [12] are inherently confined to covariate–target pairs of fixed, predefined dimension, thereby limiting their applicability to datasets that match these specifications. This limitation can be alleviated via additional data pre‑processing—e.g., projecting high‑dimensional covariates into a lower‑dimensional subspace—or by post‑processing model predictions,—e.g. employing hierarchical strategies for classification tasks with many classes [31, 5]. However, the increased computational burden often offsets some of the benefits offered by in-context learning.
@@ 2 Related Work
Modifying Prior-Fitted Networks outputs.
To the best of our knowledge, no previous approach has proposed a target-equivariant architecture for foundational tabular models, with the exception of the concurrent work from Koshil et al. [16] which unlike us avoid to process labels with non-linear outputs to focus on interpretability versus accuracy. Several works have also proposed modifications to the output architecture of TabPFN. Müller et al. [24] introduced a modification of the output, replacing the linear projection from token embeddings to the target with Generalized Additive Models. This approach improves the interpretability of the model by constructing shape functions for each feature, which can be analyzed independently. Margeloiu et al. [19] and Ye et al. [36] explored combining non-parametric models on top of TabPFN. Both approaches require training a model at inference time unlike our method.
Beyond pre-defined target dimensionality. To address TabPFN’s limitation to a predefined number of classes, Hollmann et al. [13] propose to split the classification problem into smaller ones on which the model can be employed, then to aggregate predictions using a strategy, such as the one based on an error-correcting output codes (ECOC) [5]. In this work, we show that such a strategy results in an increased computational cost compared to using our architecture that is natively target equivariant. Qu et al. [27] proposed to use a hierarchical classification strategy [31] which still incurs an increased computational cost. Recently, Wu and Bergman [35] propose a mechanism to handle arbitrary number of classes without target equivariance and require a different paradigm involving an adversarial training procedure.
Equivariance beyond tabular methods. Designing equivariant architectures [4] has long been recognized as beneficial, with the most prominent example being convolutional neural networks, which are equivariant to image translation [17]. More recently, research has focused on designing architectures with other symmetries, such as those present in spherical [9], set [37], or graph data [30]. Recent work has also explored incorporating symmetries in Large Language Models. For instance, Egressy and Stühmer [8] proposed a modification to self-attention that produces outputs equivariant to permutations of multiple-choice options, ensuring that the order of choices does not affect the result—a property known to be important in LLM applications such as LLM judges [39].
@@ 3 Background on Prior-Fitted Networks
A notable property of the TabPFN architecture is its invariance of the test predictions to the order by which training and test points are provided. This property is desirable since the order of training points is arbitrary and should not influence predictions on test samples. However, the network lacks equivariance w.r.t. the targets’ dimensions, meaning that predictions are highly dependent on the order by which the training target dimensions are provided, an undesirable property as we show in our theoretical analysis in Section˜5.
@@ 4 Target equivariant prior-fitted network
We introduce EquiTabPFN, a new model architecture for in-context learning on tabular data, that is permutation equivariant w.r.t. the target’s components. Our architecture integrates self-attention mechanisms across data points and data components to leverage relationships between datapoints while preserving equivariance by processing individual attributes (such as the targets components). Unlike TabPFN which requires fixed target dimensions for all datasets, EquiTabPFN allows the dimensions of the target to change depending on the dataset as a consequence of its equivariant architecture. EquiTabPFN consists of three major modules: a target equivariant encoder, an attention module both across data components and across data points, and a non-parametric equivariant decoder, each designed to facilitate the end-to-end learning of components interactions and datapoint relationships, see Figure˜1. Below, we elaborate on each module and their interactions.
@@ 4.1 Target equivariant encoder
The encoder constructs training and test tokens by applying a linear projection to both covariates and targets so that target equivariance is preserved. Specifically, following [12], each training and test covariate vector $x_{n}$ and $x^{\star}_{m}$ is encoded into a single token of dimension $d$ by applying a linear projection matrix $U$ of size $d\times p$. However, instead of adding a linear projection of the training targets to each corresponding training token, as done in the case of TabPFN (see Section˜3), we compute a token for each component $(y_{n})_{j}$ of a training target by multiplying them with an embedding vector $V$ of dimension $d$ for all $1\leq j\leq q$. This operation amounts to applying a $1\times 1$ convolution along the components of each target which preserves target equivariance. Since, the validation target $Y^{\star}$ is not provided to the model as input, it is replaced by a trainable prediction token $W_{\text{pred}}$ of dimension $d$ that is repeated $M\times q$ times to form an initial guess $\tilde{Y}^{0}$ of the target. When considering a batch $\mathcal{B}$ of $B$ datasets $((X,Y),(X^{\star},Y^{\star}))\in\mathcal{B}$ of same dimensions, all these embeddings along with prediction tokens are collected to form a single tensor $E$ of shape $(B,N+M,q+1,d)$. Here, for each batch element $((X,Y),(X^{\star},Y^{\star}))$ of index $b$, the blocks $E_{b,:N,1,:}$, $E_{b,N:M,1,:}$ correspond to embeddings of $X$ and $X^{\star}$, $E_{b,:N,1:q,:}$ represents the embedding of $Y$ while $E_{b,N:M,1:q,:}$ denotes the initial guess $\tilde{Y}^{0}$ obtained using the prediction token. This tensor is then processed by the attention modules as described next.
@@ 4.2 Self-Attention Mechanisms
Self-attention across components allows interactions among components within each datapoint. It is applied independently per samples to preserve equivariance w.r.t. to the samples. In practice, this is achieved by reshaping the activation into a tensor of the shape $(B\times(N+M),q+1,d)$ before applying attention between $q+1$ covariate tokens of dimension $d$. We further employ a masking strategy that we found useful empirically: forcing target tokens to attend only to the covariate token, while allowing the covariate token to attend to all tokens.
Self-Attention across datapoints captures relationships between datapoint embeddings, allowing the model to aggregate information globally. It is applied between samples and independently per each input dimensions $p$ and $q$ to preserve equivariance. In practice, this is achieved by reshaping the activation into a tensor of the shape $(B\times(q+1),N+M,d)$ before applying attention between $N+M$ training and validation tokens of dimension $d$. Similarly to [12], training and validation tokens only attend to training tokens.
Remark.
EquiTabPFN and TabPFN both have linear computational complexity in the number of classes, but EquiTabPFN has a larger factor due to self-attention across components. In TabPFN, linear scaling arises from projecting classes into a fixed-dimensional space and then mapping back fixed dimensional features to classes, while the backbone remains independent of the number of classes. In contrast, EquiTabPFN’s backbone scales linearly with class number, allowing it to handle arbitrary class counts.
@@ 4.3 Non-parametric equivariant decoder
The decoder aggregates the processed embeddings to produce prediction $\hat{Y}$. This is achieved in two steps: an attention module first computes an intermediate prediction $\tilde{Y}=(\tilde{y}_{m})_{m=1}^{M}$ in the form of a weighted average of training targets $Y$, then a residual correction is added to produce the final prediction. More precisely, the attention module uses the embeddings of the training and validation samples as keys and queries, while the attention values are simply the training targets $Y$, i.e. $\tilde{y}_{m}=\sum_{n=1}^{N}y_{n}\text{SoftMax}\left(\sum_{i,u}E_{b,n,i,u}E_{b,m,i,u}/\sqrt{(1+q)d}\right),$ where $b$ is the batch-index corresponding to the training target $Y$. The residual correction, in the form of a point-wise MLP, operates independently on each dimension $j$ of the attention output $\tilde{y}_{m}$ so that equivariance is preserved while enabling nonlinear interactions between training values.
Without the residual correction and removing the dependence of the keys and queries embeddings on the training targets $Y$ (for instance by setting the weights of the target encoder and pointwise MLP to 0), the decoder becomes a linear non-parametric regression estimator [32, Definition 1.7], which is a generalization of Nadaraya-Watson’s estimator [25, 34]. However, linear estimators are known to be suboptimal compared to non-linear ones [6]. This motivates introducing a nonlinear dependence of the estimator to $Y$, in our setting, to increase the expressiveness of the decoder allowing it to adapt to the prediction task at hand. Experimentally, we also found a clear improvement when adding such a residual correction and making the embeddings dependent on the targets.
@@ 4.4 Pre-training Procedure
Here, $(y,y^{\prime})\mapsto\ell(y,y^{\prime})\in\mathbb{R}$ is a point-wise loss, typically cross-entropy, and the expectation is over the collections datasets sampled according to the dataset prior. Note that each test query $x_{m}^{\star}$ is processed independently by the network $f$ so that $f_{X,Y}\left(X^{\star}\right)=(f_{X,Y}\left(x_{m}^{\star}\right))_{m=1}^{M}$. Next, we show, under natural conditions, that target equivariant functions constitute the right space of functions when searching for solutions to objectives of the form in Equation˜1.
@@ 5.1 Optimality of target equivariant networks
It is clear from Definition˜5.1 that EquiTabPFN is target equivariant. Even when a model is not target equivariant by construction, it is natural to expect it to learn to be target equivariant, when trained via the objective in Equation˜1 over a large class of randomly sampled datasets. To formalize this, we define the target equivariance gap, which quantifies the deviation of a function $f$ from its symmetrized counterpart ${f}^{\text{equi}}$.
@@
Assumption˜(A) is satisfied for most commonly used losses such as the cross-entropy loss or the quadratic loss. Assumption˜(B) holds as soon as data can be presented without a preferred ordering, as most i.i.d. tabular datasets. The next proposition, proved in Appendix˜A, decomposes the objective into a non-negative equivariance gap and an optimality error.
@@
Under Assumptions˜(A) and (B), the equivariance gap $\mathcal{E}^{\text{equi}}(f)$ is always non-negative and only equal to 0 when $f$ is equivariant to permutations, so that for any $f$:
@@
The key step is to express the objective as an expectation over permutations using Assumptions˜(A) and (B) on the data and loss: $\mathcal{L}(f)=\mathbb{E}_{p}\mathbb{E}_{\sigma}\left[\ell\left(\sigma^{-1}f_{X,\sigma(Y)}\left(X^{\star}\right),Y^{\star}\right)\right]$. The non-negativity of the equivariance gap is then established using Jensen’s inequality by convexity of the loss $\ell$ in its first argument (Assumption˜(A)). Now, assume by contradiction that $f^{\star}$ is not equivariant and note that ${f^{\star}}^{\text{equi}}$ is a measurable function by construction. It follows that $\mathcal{E}^{\text{equi}}(f^{\star})>0$, which directly implies that $\mathcal{L}(f^{\star})>\mathcal{L}({f^{\star}}^{\text{equi}})$, thus contradicting the optimality of $f^{\star}$. ∎
Proposition˜5.3 shows that minimizing the objective $\mathcal{L}$ results in a target equivariant function. Hence, using a non-equivariant model must employ some of its expressive power solely for being equivariant, which can be wasteful, as we verify empirically in Section˜5.2 below.
@@ 5.2 Non-equivariance of TabPFNs models
Predictions instabilities. To illustrate the implications of non-equivariance on the robustness of PFNs models, we consider a toy classification problem in $2$ dimensions, where $9$ training points are positioned on a $2$-dimensional regular grid, each corresponding to a different class inspired by McCarter [20]. Different pretrained models are then used to predict the classes on a regular grid of $40^{2}$ points. Figure˜3 shows the classification boundaries when using the same data but with $3$ different orderings for the classes. It is clear that the ordering heavily affects the prediction even in this simple example, which strongly impacts robustness of models like TabPFN and TabPFN-v2. The predictions of TabPFN-v2 are particularly noisy due to having only 9 training data points which is not an issue in itself, unlike the extreme unstability to the class ordering. Note that the axis are scaled differently for presentation.
Target equivariance gap during training. In Figure˜3, we analyse the equivariance error while training TabPFN. Figure˜3 (left) shows the equivariance error of TabPFN in terms of percentage of violation of Equation˜2, e.g. how frequently the predicted classes $f_{X,\sigma(Y)}(X^{*})$ and $\sigma\left(f_{X,Y}(X^{*})\right)$ differ. We sample 512 datasets from the prior and report standard deviation. In Figure˜3 (left), the equivariance error is clear and slowly decreases during training. This non-equivariance 1) induces additional errors for the model as demonstrated in Proposition˜5.3 and 2) causes the model to provide surprising predictions given that permuting the output order can change the results as seen in Figure˜3.
Mitigation via costly ensembling. Hollmann et al. [12] introduce a mitigation strategy to non-equivariance by averaging several predictions using random permutations of the target dimensions. Averaging over all possible permutations gives an equivariant function as discussed in Equation˜4. However, this requires making $\mathcal{O}(q!)$ calls to the model, where $q$ is the number of classes. This becomes quickly prohibitive, even for $q=10$ as considered in the original study. Randomized estimators, using $N_{ens}$ random permutations can compute a prediction that converge to the averaged one at a rate of $1/\sqrt{N_{ens}}$. However, the variance of such estimator would typically present a dependence on the dimension that is, unfortunately, challenging to quantify in general. This variance is highest when the function is far from being equivariant, constituting the worst scenarios. We illustrate how fast the model becomes equivariant when using ensembles in Figure˜3 (right). While ensembling more permutation helps to make the model more equivariant, many ensembles are required, in particular, when considering more classes. In contrast, the model we propose is fully target equivariant so that predictions remain the same regardless of the class ordering as illustrated in Figure˜3 (bottom).
@@ 6.1 Experimental setup
We trained our model, EquiTabPFN on artificial dataset extending the public code of Müller et al. [22] and using the same procedure from Hollmann et al. [12] employed to train TabPFNv1. Crucially, it was trained on classification tasks with less than $10$ classes. The total training time took approximately 4 days on a single A100 GPU with 80GB of memory. We set our architectures hyperparameters so that they match the number of parameters of previous work, in particular the model we trained contains $\sim$25M parameters similar to the baselines we consider and we match most our hyperparameters to the same values as Hollmann et al. [12]. We refer to Appendix˜B for further training details and description of the hyperparameters used.
@@
For evaluation, we consider classification tasks from the TabZilla benchmark [21]. To assess the impact of unseen class counts during evaluation (i.e., exceeding $10$), we consider two setups: one on $76$ multi-class tasks with at most $10$ classes, and another on $10$ tasks with more than $10$ classes. Details of the datasets/tasks used are provided in Tables˜1 and 2. In all results, except those of Figure˜5 (left), we report the results of EquiTabPFN without ensembling as it is not critical for the performance of our method.
@@
6.2 Main results
This section presents the main experimental findings. Additional results, including ablation of the EquiTabPFN architecture and computational cost comparisons are found in Sections˜C.1 and C.2.
Figure 5: Left: Scatter plot of runtime vs $\%$ AUC improvement over KNN for different methods. Right: barplot of AUC ratio relatively to KNN (blue) and speedup of EquiTabPFN (green). In both figures, we increase the number of ensembles of TabPFNs variants.